OpenStudy (anonymous):

You are given the rate of rotation of a wheel as well as its radius. Determine the following. (Express your answers in terms of π.) 1210 rpm; r = 28 cm (a) angular speed, in units of radians/sec B)linear speed, in cm/sec, of a point halfway between the center of the wheel and the circumference

OpenStudy (anonymous):

a) 7*2π = 43.98 rad/s b) v = r* ω = 11*43.98 = 483.8 cm/s its right off of google but its understandable. <3